OpenStudy (anonymous):

x + y = 6 x - y = 8 Solve the system of equations. i got (1, -7)?

OpenStudy (anonymous):

x = 16/7, y = 40/7 but i think im wrong

OpenStudy (anonymous):

\(x+y = 6\) + \(x -y = 8\) =========== \(2x + 0y = 14\) \[\implies x = 7\]

OpenStudy (anonymous):

From there you can find y by plugging in that value of x into either of the original equations

OpenStudy (anonymous):

right